START-UP DEMONSTRATION TESTS UNDER

MARKOV DEPENDENCE MODEL WITH CORRECTIVE ACTIONS

N. BALAKRISHNAN^{1}, S. G. MOHANTY^{1} AND S. AKI^{2}

1Department of Mathematics and Statistics, McMaster University, Hamilton, Ontario, Canada L8S 4K1

2Department of Mathematical Science, Osaka University, Toyonaka, Osaka 560, Japan

(Received January 22, 1996; revised May 20, 1996)

Abstract. A general probability model for a start-up demonstration test is studied. The joint probability generating function of some random variables appearing in the Markov dependence model of the start-up demonstration test with corrective actions is derived by the method of probability generating function. By using the probability generating function, several characteristics relating to the distribution are obtained.

Key words and phrases: Probability generating function, discrete distributions, run, start-up demonstration test, Markov chain, waiting time.